Python 3

Write a binary search function that, instead of returning -1 when the target value is not in the list, raises a TargetNotFound exception (you'll need to define this exception class). Otherwise it should function normally. Name this function bin_except.

The file must be named: bin_except.py
Answer #1

- targetnotfound.py - D:/Coding/Python/targetnotfound.py (3.7... File Edit Format Run Options Window Help #Define TargetNot F

OUTPUT :

================ RESTART: D: /Coding/Python/targetnotfound.py >>> a = [1,2,5,9,15] >>> bin_except (a, 10) Traceback (most rec

CODE :

#Define TargetNotFound exception
class TargetNotFound(Exception):
pass

#Function that performs binary search
def bin_except(a,key):
#Initialize l and r
l = 0
r = len(a)-1
#Continue till l<=r
while l<=r:
#Find mid value m
m = (l+r)//2
#Check if key is found
if a[m]==key:
return m
#Change l or r
if a[m]<key:
l = m+1
else:
r = m-1
raise TargetNotFound